OpenStudy (ray_sihota):

Help Health food magazine has 14 advertising sales people. 4 men and 10 women. 6 salespeople will be randomly selected to attend the annual health food convention in Las Vegas. What is the probability at of at least 3 men being selected.

jimthompson5910 (jim_thompson5910):

`at least 3 men` means 3 or more men. There are 4 men total, so we could have 3 men or 4 men selected. There are no other possible options. So what I would do is break up the problem into 2 cases Case A will be where 3 men are selected. Case B will be where 4 men are selected. ----------------------------------------- Case A) 3 men are selected Use the combination formula to determine how many ways we can pick 3 men \[\Large _n C _r = \frac{n!}{r!*(n-r)!}\] \[\Large _4 C _3 = \frac{4!}{3!*(4-3)!}\] \[\Large _4 C _3 = \frac{4!}{3!*1!}\] \[\Large _4 C _3 = \frac{4*3*2*1}{3*2*1*1}\] \[\Large _4 C _3 = \frac{24}{6}\] \[\Large _4 C _3 = 4\] There are 4 ways to select 3 men from a pool of 4. Use the same formula to determine how many ways there are to pick 3 women \[\Large _n C _r = \frac{n!}{r!*(n-r)!}\] \[\Large _{10} C _3 = \frac{10!}{3!*(10-3)!}\] \[\Large _{10} C _3 = \frac{10!}{3!*7!}\] \[\Large _{10} C _3 = \frac{10*9*8*7*6*5*4*3*2*1}{3*2*1*7*6*5*4*3*2*1}\] \[\Large _{10} C _3 = \frac{3,628,800}{30,240}\] \[\Large _{10} C _3 = 120\] There are 120 ways to select 3 women from a pool of 10 Multiply the previous results 4*120 = 480 So there are 480 ways to pick 3 men and 3 women (from a pool of 4 men and 10 women) Hopefully this makes sense?

jimthompson5910 (jim_thompson5910):

let me know if you're able to do case B @ray_sihota

OpenStudy (ray_sihota):

than you it does make sense for case b I did 4c4 x10 c 2 and added that to 480 which give me 525/14 c 6 525/3003

jimthompson5910 (jim_thompson5910):

You should find that 4 C 4 = 1 10 C 2 = 45 so 1*45 = 45 There are 45 ways to pick 4 men and 2 women (pool of 4 men and 10 women) add this to the previous result from case A to get 480+45 = 525 so you have the correct values. Nice work

jimthompson5910 (jim_thompson5910):

There are 525 ways to either pick 3 men or pick 4 men. This value will be the numerator to be used later Let's calculate the denominator There are 4+10 = 14 people. There are 6 slots to fill. What is the value of 14 C 6 ?

jimthompson5910 (jim_thompson5910):

oh you already wrote it out, you got 3003 which is correct so the probability is 525/3003

jimthompson5910 (jim_thompson5910):

make sure you reduce the fraction as much as possible. In this case, 525/3003 = 25/143 so 25/143 is the final answer assuming your teacher wants a fraction as the answer. If s/he wants a decimal, then use a calculator to get the decimal form of 25/143